OpenStudy (anonymous):

I need help to solve this problem 5w-3(6.4)=120+0.2(38) w=29.36 but I cant find the solution to solve it

OpenStudy (amistre64):

this one right? just checking it out..

OpenStudy (amistre64):

5w -3(6.4) = 120+ 0.2(38) 5w = 120+ 0.2(38) +3(6.4) w = 120+ 0.2(38) +3(6.4) ------------------ 5

OpenStudy (anonymous):

5w-19.2=120+7.6 5w-19.2=127.6 5w=146.8 w=29.36

OpenStudy (amistre64):

w =29.36

OpenStudy (anonymous):

not this one...difference equations for me

OpenStudy (anonymous):

multiply on both sides add on right side add 19.2 to both sides divide by 5